500ml of a 0.1N solution of AgNO3 added to 500 ml of 0.1 N solution of KCl.The concentration of Nitrate Ion in the resulting solution is what?

Grade:11

1 Answers

Arun
4 years ago
 
Gram equivalents of NO3 in AgNO3 = 0.500 × 0.1 = 0.05 gmeq
There is no gmeqs of NO3 in KCl.
Total volume = 500ml + 500ml = 1L
Hence, concentration of NO3 = 0.05 / 1 = 0.05N